OpenStudy (anonymous):

1.1. Is the equation true, false, or open? Explain. 9p + 8 = 10p + 7 2. Is x = 1 a solution of the equation 2 – 8x = –6? 3. A souvenir maker wants to create a scale model of the Eiffel Tower. The Eiffel Tower is 324 meters tall and has a base with dimensions 125 meters by 125 meters. The model will rest on a base with dimensions 12.5 cm by 12.5 cm. How tall will the model be in centimeters?

OpenStudy (anonymous):

If you solve for p, you will get p=1. That means it is true. If p could equal a whole range of values, then it would be open. If you get something like nonsensical like 1=2 when you try to solve, it then it would be false.... 2-8x=-6 -2 -8 -8x=-8 -8=1 Yes 125/12.5 = 10 so the scale is 10cm:1m 324/10 = 32.4cm